OpenStudy (anonymous):

Someone help!!! X=-3y^2+30y Please explain how to do the equation step by step so i can understand.

OpenStudy (anonymous):

what does "do \(x=-3y^2+3y\)" mean?

OpenStudy (anonymous):

possibly factor it

OpenStudy (anonymous):

maybe it means find the parabola, vertex, focus etc but maybe not are there any instructions?

OpenStudy (anonymous):

there is no such mathematical operation as "do"

OpenStudy (anonymous):

Hey, my computer shut down and I had to go through a lot to turn it on. Sorry :(

OpenStudy (anonymous):

So I know this is a parabola, its just I'm having a tough time trying to put it in standard form

OpenStudy (anonymous):

oooh ok

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

\[x=-3y^2+30y \]\[x=-3(y^2-10y)\] \[x=-3(y-5)^2+75\]

OpenStudy (anonymous):

or \[x-75=-3(y-5)^2\]

OpenStudy (anonymous):

How did you get 75

OpenStudy (anonymous):

good question did you get the first step?

OpenStudy (anonymous):

Yes, -3(y^2-10y), you factored out the -3,

OpenStudy (anonymous):

And then you divided the 10 by 2,

OpenStudy (anonymous):

then to complete the square take have of ten which is 5 and write \[x=-3(y-5)^3+??\]

OpenStudy (anonymous):

25+?

OpenStudy (anonymous):

I don't know the other number

OpenStudy (anonymous):

now \(5^2=25\) and \(-3\times 25=-75\) so by making that change you have subtracted 75 you can either subtract it from the other side or add it back to the right side

OpenStudy (anonymous):

thanks, pal Now my brain woke up. That's the last step I was missing.

OpenStudy (anonymous):

I'm blast234

OpenStudy (anonymous):

I have sibilings who come here :D

OpenStudy (anonymous):

must be a busy computer

OpenStudy (anonymous):

Maybe thats why, but most of them are unfamiliar